OpenStudy (anonymous):

Which of the following statements are true of the given lines? Line A: 2x+y=-4 Line B: X+ 2y= -10 Line C: -2x+ 4y= -12 a. Lines a & b are parallel b. Lines b & c are parallel c. Lines a & c are perpendicular d. Lines b & c are perpendicular

OpenStudy (anonymous):

ok first set all of the equations like, y=mx+b

OpenStudy (anonymous):

Line A: y= -2x-4 Line B: y= -0.5x-5 Line C: y= 0.5x-3

OpenStudy (anonymous):

yass

OpenStudy (anonymous):

ok so what would parallel n perpendicular mean?

OpenStudy (anonymous):

Will Lines B&C be parallel?

OpenStudy (anonymous):

well parallel would be the exact slope

OpenStudy (anonymous):

So they are a perpendicular?

OpenStudy (anonymous):

no perpendicular means opposite in sign n its flipped

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

ok perpendicular is, 6/5 and -5/6

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

Yes, I get that but the problem is where I am confused

OpenStudy (anonymous):

ok you have, -1/2 and 2 and 1/2, which ones are perpendicular?

OpenStudy (anonymous):

2 and 1/2

OpenStudy (anonymous):

no it would be 2 and -1/2 or -2 and 1/2, get it?

OpenStudy (anonymous):

ohhh okay now I get it

OpenStudy (anonymous):

so whats ur answer now, for you problem?

OpenStudy (anonymous):

letter c

OpenStudy (anonymous):

a and c are perpendicular

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you welcome cx

OpenStudy (anonymous):

thank you, you know much about math

OpenStudy (anonymous):

im a math nerd xp